### v1

# Toric intersections

Abstract.
Let $I_{A}\subset K[x_{1},\ldots,x_{n}]$ be a toric ideal. In this paper, we provide a necessary and sufficient condition for the toric variety $V(I_{A})$, over an algebraically closed field, to be expressed as the set-theoretic intersection of other toric varieties. We also introduce the invariant ${\rm Split}_{\rm rad}(I_{A})$, defined as the smallest integer $r$ for which there exist toric ideals $I_{A_{1}},\dots,I_{A_{r}}$ satisfying $I_{A}={\rm rad}(I_{A_{1}}+\cdots+I_{A_{r}})$ and $I_{A_{i}}\neq I_{A}$ for all $1\leq i\leq r$. We then compute its exact value in several cases, including the case in which $I_{A}$ is the toric ideal of a complete bipartite graph. Additionally, we show that ${\rm Split}_{\rm rad}(I_{A})$ is equal to the binomial arithmetical rank of $I_{A}$ when the height of $I_{A}$ is equal to 2.
Key words and phrases:
Radical splittings of toric ideals, minimal systems of binomial generators up to radical, binomial arithmetical rank, bipartite graphs.
1991 Mathematics Subject Classification:
14M25, 13F65, 05C25, 05E40.
Corresponding author: Anargyros Katsabekis
1. Introduction
A natural question in algebraic geometry is to determine when an affine algebraic variety $V\subset K^{m}$ can be expressed as the intersection of two distinct affine algebraic varieties, $V_{1}$ and $V_{2}$, over an algebraically closed field $K$; that is, $V=V_{1}\cap V_{2}$, with $V\subsetneqq V_{i}$ for $1\leq i\leq 2$. This geometric question translates algebraically via the Hilbert Nullstellensatz into an ideal-theoretic question:
$I(V)={\rm rad}(I(V_{1})+I(V_{2})),$
where $I(W)$ denotes the ideal of an affine algebraic variety $W$ and ${\rm rad}(J)$ is the radical of an ideal $J$, see [4, Chapter 4]. This paper addresses the latter problem in the case where all varieties involved are toric varieties. In other words, we study when a given toric ideal $I_{A}\subset K[x_{1},\ldots,x_{n}]$ can be written as $I_{A}={\rm rad}(I_{A_{1}}+I_{A_{2}})$ and $I_{A_{i}}\subsetneqq I_{A}$ for $1\leq i\leq 2$. Recent advances in understanding toric splittings [6, 7, 11, 12], along with several earlier results on the binomial generation of toric ideals up to radical [8, 9], allow us to provide a necessary and sufficient condition for the equality $I_{A}={\rm rad}(I_{A_{1}}+I_{A_{2}})$.
Let $A=\{{\bf a}_{1},\ldots,{\bf a}_{n}\}$ be a vector configuration in $\mathbb{Z}^{m}$, ${\rm ker}_{\mathbb{Q}}(A)=\{{\bf u}=(u_{1},\ldots,u_{n})\in\mathbb{Q}^{n}\mid u_{1}{\bf a}_{1}+\cdots+u_{n}{\bf a}_{n}={\bf 0}\}$, and ${\rm ker}_{\mathbb{Z}}(A)={\rm ker}_{\mathbb{Q}}(A)\cap\mathbb{Z}^{n}$. Throughout this paper, we will assume that the affine semigroup $\mathbb{N}A=\{l_{1}{\bf a}_{1}+\cdots+l_{n}{\bf a}_{n}\mid l_{i}\in\mathbb{N}\}$ is pointed; namely ${\rm ker}_{\mathbb{Z}}(A)\cap\mathbb{N}^{n}=\{{\bf 0}\}$. Consider the polynomial ring $K[x_{1},\ldots,x_{n}]$ over a field $K$. The toric ideal $I_{A}$ is the kernel of the $K$-algebra homomorphism $\phi:K[x_{1},\ldots,x_{n}]\longrightarrow K[t_{1},\ldots,t_{m},t_{1}^{-1},\ldots,t_{m}^{-1}]$ given by $\phi(x_{i})={\bf t}^{{\bf a}_{i}}$ for all $i=1,\ldots,n$ and it is generated by all the binomials $B({\bf u}):={\bf x}^{{\bf u}^{+}}-{\bf x}^{{\bf u}^{-}}$ such that ${\bf u}^{+}-{\bf u}^{-}\in{\rm ker}_{\mathbb{Z}}(A)$, where ${\bf u}^{+}\in\mathbb{N}^{n}$ and ${\bf u}^{-}\in\mathbb{N}^{n}$ denote the positive and negative part of ${\bf u}={\bf u}^{+}-{\bf u}^{-}$, respectively (see, for instance, [16, Lemma 4.1]). By [16, Lemma 4.2], the height ${\rm ht}(I_{A})$ of $I_{A}$ equals the dimension of the vector space ${\rm ker}_{\mathbb{Q}}(A)$. The set $V(I_{A})=\{(u_{1},\ldots,u_{m})\in K^{m}\mid F(u_{1},\ldots,u_{m})=0,\forall F\in I_{A}\}$ of zeroes of $I_{A}$ is called affine toric variety, which is not necessarily normal.
The toric ideal $I_{A}$ is splittable if there exist toric ideals $I_{A_{1}},I_{A_{2}}$ such that $I_{A}=I_{A_{1}}+I_{A_{2}}$ and $I_{A_{i}}\not=I_{A}$ for all $1\leq i\leq 2$. The ideal $I_{A}$ is called radical splittable if and only if there exist toric ideals $I_{A_{1}},I_{A_{2}}$ such that $I_{A}={\rm rad}(I_{A_{1}}+I_{A_{2}})$ and $I_{A_{i}}\neq I_{A}$, for $1\leq i\leq 2$.
Definition 1.1.
Let $I_{A}$ be a toric ideal.
  1. (1)
    The splitting number of $I_{A}$, denoted by ${\rm Split}(I_{A})$, is the smallest integer $s$ such that $I_{A}=I_{A_{1}}+\cdots+I_{A_{s}}$ and $I_{A_{i}}\neq I_{A}$ for every $1\leq i\leq s$. In this case, $I_{A}=I_{A_{1}}+\cdots+I_{A_{s}}$ is a splitting of $I_{A}$.
  2. (2)
    The radical splitting number of $I_{A}$, denoted by ${\rm Split}_{\mathrm{rad}}(I_{A})$, is the smallest integer $r$ such that $I_{A}={\rm rad}(I_{A_{1}}+\cdots+I_{A_{r}})$ and $I_{A_{i}}\neq I_{A}$ for every $1\leq i\leq r$. In this case, $I_{A}={\rm rad}(I_{A_{1}}+\cdots+I_{A_{r}})$ is a radical splitting of $I_{A}$.
Remark 1.2.
(1) Since $I_{A}$ is minimally generated by binomials, we get that ${\rm Split}(I_{A})\leq\mu(I_{A})$, where $\mu(I_{A})$ is the minimal number of generators of $I_{A}$.
(2) The invariant ${\rm Split}(I_{A})$ is an upper bound for ${\rm Split}_{\mathrm{rad}}(I_{A})$.
The binomial arithmetical rank of $I_{A}$, denoted by ${\rm bar}(I_{A})$, is the smallest integer $t$ for which there exist binomials $B_{1},\ldots,B_{t}$ in $I_{A}$ such that $I_{A}={\rm rad}(B_{1},\ldots,B_{t})$. From the generalized Krull’s principal ideal theorem, we deduce that ${\rm ht}(I_{A})$ is a lower bound for ${\rm bar}(I_{A})$. From the definitions, we obtain the following inequalities for a non-principal toric ideal $I_{A}$:
$2\leq{\rm Split}_{\mathrm{rad}}(I_{A})\leq{\rm bar}(I_{A})\leq\mu(I_{A}).$
For a non-principal toric ideal $I_{A}$ we have that ${\rm Split}_{\mathrm{rad}}(I_{A})=2$ if and only if $I_{A}$ is radical splittable.
In Section 2, we provide a necessary and sufficient condition for a toric ideal $I_{A}$ to be radical splittable (see Corollary 2.2). Also, we show that if the height of $I_{A}$ is $r\geq 3$ and ${\rm bar}(I_{A})\leq 2r-2$, then $I_{A}$ is radical splittable (see Theorem 2.5). Moreover, if $I_{A}$ is a toric ideal of height $r\geq 3$ such that $V(I_{A})$ is a simplicial toric variety with full parametrization, then $I_{A}$ is radical splittable (see Proposition 2.8).
In Section 3, we prove that for a bipartite graph $G$, ${\rm Split}(I_{G})={\rm Split}_{\rm rad}(I_{G})$ holds, where $I_{G}$ is the toric ideal of the graph $G$ (see Theorem 3.1). Furthermore, we show that ${\rm Split}_{\rm rad}(I_{G})=3$ for a complete bipartite graph $G$ (see Theorem 3.3).
In Section 4, we study the special case in which $I_{A}$ is a toric ideal of height 2. We show that ${\rm Split}(I_{A})=\mu(I_{A})$ and ${\rm Split}_{\rm rad}(I_{A})={\rm bar}(I_{A})$ (see Theorem 4.1). We also show (see Corollary 4.4) that if $V(I_{A})$ is a simplicial toric variety, then $2\leq{\rm Split}_{\rm rad}(I_{A})\leq 3$, and both lower and upper bounds are attained; see Example 4.5. Furthermore, we explicitly compute the splitting and the radical splitting number of the toric ideal $I_{\Lambda(A)}$ (see Proposition 4.8), where $\Lambda(A)$ is the Lawrence lifting of the vector configuration $A$ associated with a symmetric monomial curve in $\mathbb{P}^{3}$. Proposition 4.8 shows that the splitting number can be arbitrarily large for some configurations, while the radical splitting number remains constant for them. Finally, we show that both ${\rm Split}(I_{A})$ and ${\rm Split}_{\rm rad}(I_{A})$ can be arbitrarily large for certain other configurations; see Proposition 4.9.
2. Radical Splitting Criterion for $I_{A}$
Let $A=\{{\bf a}_{1},\ldots,{\bf a}_{n}\}$ be a vector configuration in $\mathbb{Z}^{m}$. Given a set $C=\{{\bf c}_{1},\ldots,{\bf c}_{k}\}\subset\mathbb{Z}^{n}$, we shall denote by ${\rm span}_{\mathbb{Q}}(C)$ the $\mathbb{Q}$-vector space generated by the vectors of $C$, namely,
${\rm span}_{\mathbb{Q}}(C)=\{\lambda_{1}{\bf c}_{1}+\cdots+\lambda_{k}{\bf c}_{k}\mid\lambda_{i}\in\mathbb{Q}\}.$
A set $S$ is a minimal system of binomial generators of the toric ideal $I_{A}$ up to radical if there exist no $S^{\prime}\subsetneqq S$ such that $S^{\prime}$ generate $I_{A}$ up to radical.
Theorem 2.1.
There are toric ideals $I_{A_{i}}$ with $I_{A_{i}}\neq I_{A}$ for every $1\leq i\leq s$, such that $I_{A}=I_{A_{1}}+\cdots+I_{A_{s}}$ (respectively, $I_{A}={\rm rad}(I_{A_{1}}+\cdots+I_{A_{s}})$) if and only if there exists a minimal system of binomial generators $\{B({\bf u})\mid{\bf u}\in C\subset{\rm ker}_{\mathbb{Z}}(A)\}$ of the toric ideal $I_{A}$ (respectively, of $I_{A}$ up to radical), and sets $C_{i}$, $1\leq i\leq s$, such that $C=\cup_{i=1}^{s}C_{i}$ and ${\rm span}_{\mathbb{Q}}(C_{i})\subsetneqq{\rm ker}_{\mathbb{Q}}(A)$ for every $1\leq i\leq s$.
Proof. Suppose that $I_{A}=I_{A_{1}}+\cdots+I_{A_{s}}$ (respectively, $I_{A}={\rm rad}(I_{A_{1}}+\cdots+I_{A_{s}})$) and $I_{A_{i}}\neq I_{A}$, for every $1\leq i\leq s$. Thus, ${\rm ker}_{\mathbb{Z}}(A_{i})\subsetneqq{\rm ker}_{\mathbb{Z}}(A)$; therefore, ${\rm ker}_{\mathbb{Q}}(A_{i})\subsetneqq{\rm ker}_{\mathbb{Q}}(A)$ for every $1\leq i\leq s$. Let $T_{i}$, $1\leq i\leq s$, be a minimal system of binomial generators of $I_{A_{i}}$ (respectively, of $I_{A_{i}}$ up to radical). Then, since $I_{A}=I_{A_{1}}+\cdots+I_{A_{s}}$ (respectively, $I_{A}={\rm rad}(I_{A_{1}}+\cdots+I_{A_{s}})$), we have that $D=\cup_{i=1}^{s}T_{i}$ generates $I_{A}$ (respectively, $I_{A}$ up to radical). Thus, there exists a minimal system of binomial generators of $I_{A}$ (respectively, of $I_{A}$ up to radical) in the form $\{B({\bf u})\mid{\bf u}\in C\subset D\}$. Let $C_{i}=D\cap T_{i}$ for $1\leq i\leq s$. Thus, $C=\cup_{i=1}^{s}C_{i}$. Since $C_{i}\subset{\rm ker}_{\mathbb{Z}}(A_{i})$, we have ${\rm span}_{\mathbb{Q}}(C_{i})\subset{\rm ker}_{\mathbb{Q}}(A_{i})$; therefore, ${\rm span}_{\mathbb{Q}}(C_{i})\subsetneqq{\rm ker}_{\mathbb{Q}}(A)$ for every $1\leq i\leq s$.
Conversely, suppose that there exists a minimal system of binomial generators $\{B({\bf u}_{1}),\ldots,B({\bf u}_{s})\mid{\bf u}_{i}\in C,1\leq i\leq s\}$ of the toric ideal $I_{A}$ (respectively, of $I_{A}$ up to radical), where $C\subset{\rm ker}_{\mathbb{Z}}(A)$, and sets $C_{i}$, $1\leq i\leq s$, such that $C=\cup_{i=1}^{s}C_{i}$ and ${\rm span}_{\mathbb{Q}}(C_{i})\subsetneqq{\rm ker}_{\mathbb{Q}}(A)$ for every $1\leq i\leq s$. The set ${\rm span}_{\mathbb{Q}}(C_{i})$ is a subspace of $\mathbb{Q}^{n}$. For every $1\leq i\leq s$, consider the matrix $M_{i}$ whose rows form a generating set for the orthogonal complement of the vector space ${\rm span}_{\mathbb{Q}}(C_{i})$. Let $A_{i}$ be the set of columns of $M_{i}$. Since $({\rm span}_{\mathbb{Q}}(C_{i}))^{\bot\bot}={\rm ker}_{\mathbb{Q}}(A_{i})$ and also $({\rm span}_{\mathbb{Q}}(C_{i}))^{\bot\bot}={\rm span}_{\mathbb{Q}}(C_{i})$, it follows that ${\rm ker}_{\mathbb{Q}}(A_{i})={\rm span}_{\mathbb{Q}}(C_{i})$ for every $1\leq i\leq s$.
We claim that $I_{A}=I_{A_{1}}+\cdots+I_{A_{s}}$ (respectively, $I_{A}={\rm rad}(I_{A_{1}}+\cdots+I_{A_{s}})$) is a splitting of $I_{A}$ (respectively, radical splitting of $I_{A}$). Since $C_{i}\subset{\rm ker}_{\mathbb{Z}}(A)$, it follows that ${\rm ker}_{\mathbb{Z}}(A_{i})\subset{\rm ker}_{\mathbb{Z}}(A)$. Therefore, $I_{A_{i}}\subset I_{A}$ for every $1\leq i\leq s$, and thus ${\rm rad}(I_{A_{1}}+\cdots+I_{A_{s}})\subset I_{A}$. For each generator $B({\bf u}_{i})$, where $1\leq i\leq s$, of $I_{A}$ (respectively, of $I_{A}$ up to radical), we have that ${\bf u}_{i}\in C_{j}$ for $1\leq j\leq s$. Thus $B({\bf u}_{i})\in I_{A_{j}}$. Therefore, every $B({\bf u}_{i})$, where $1\leq i\leq s$, belongs to at least one of $I_{A_{1}},\ldots,I_{A_{s}}$. We conclude that $I_{A}\subset I_{A_{1}}+\cdots+I_{A_{s}}$ (respectively, $I_{A}\subset{\rm rad}(I_{A_{1}}+\cdots+I_{A_{s}})$), and thus $I_{A}=I_{A_{1}}+\cdots+I_{A_{s}}$ (respectively, $I_{A}={\rm rad}(I_{A_{1}}+\cdots+I_{A_{s}})$).
Note also that $I_{A_{i}}\not=I_{A}$ for $1\leq i\leq s$. Otherwise, if $I_{A_{i}}=I_{A}$, then ${\rm ker}_{\mathbb{Z}}(A)={\rm ker}_{\mathbb{Z}}(A_{i})$, and therefore ${\rm ker}_{\mathbb{Q}}(A)={\rm span}_{\mathbb{Q}}(C_{i})$, which is a contradiction. $\square$
The next corollary is derived directly from Theorem 2.1 for $s=2$.
Corollary 2.2.
The toric ideal $I_{A}$ is radical splittable if and only if there exists a minimal system of binomial generators $\{B({\bf u})\mid{\bf u}\in C\subset{\rm ker}_{\mathbb{Z}}(A)\}$ of the toric ideal $I_{A}$ up to radical, and sets $C_{1}$ and $C_{2}$ such that $C=C_{1}\cup C_{2}$, ${\rm span}_{\mathbb{Q}}(C_{1})\subsetneqq{\rm ker}_{\mathbb{Q}}(A)$, and ${\rm span}_{\mathbb{Q}}(C_{2})\subsetneqq{\rm ker}_{\mathbb{Q}}(A)$.
The following proposition establishes that $I_{A}$ is radical splittable whenever it is a set-theoretic complete intersection on binomials; that is, when its binomial arithmetical rank coincides with its height.
Proposition 2.3.
Let $K$ be a field of any characteristic and let $I_{A}$ be a toric ideal of height $r\geq 2$. If $I_{A}$ is a set-theoretic complete intersection on binomials, then $I_{A}$ is radical splittable.
Proof. Suppose that $I_{A}={\rm rad}(B({\bf u}_{1}),\ldots,B({\bf u}_{r}))$. Let $C=\{{\bf u}_{1},\ldots,{\bf u}_{r}\}$, $C_{1}=\{{\bf u}_{1}\}$, and $C_{2}=\{{\bf u}_{2},\ldots,{\bf u}_{r}\}$. Then ${\rm span}_{\mathbb{Q}}(C_{1})\subsetneqq{\rm ker}_{\mathbb{Q}}(A)$ and ${\rm span}_{\mathbb{Q}}(C_{2})\subsetneqq{\rm ker}_{\mathbb{Q}}(A)$, since the dimension of ${\rm ker}_{\mathbb{Q}}(A)$ is equal to $r$ and the cardinality of $C_{i}$, $1\leq i\leq 2$, is strictly less than $r$. By Corollary 2.2, the toric ideal $I_{A}$ is radical splittable. $\square$
Remark 2.4.
Note that Proposition 2.3 is characteristic free, but the property of $I_{A}$ being a set-theoretic complete intersection on binomials depends on the characteristic. For details, see [2].
The next theorem shows that the corresponding statement of Proposition 2.3 also holds when ${\rm bar}(I_{A})\leq 2{\rm ht}(I_{A})-2$, provided that ${\rm ht}(I_{A})\geq 3$.
Theorem 2.5.
If $I_{A}$ is a toric ideal of height $r\geq 3$ such that its binomial arithmetical rank is less than or equal to $22r-$ 2, then $I_{A}$ is radical splittable.
Proof. Let $\{B({\bf u}_{1}),\ldots,B({\bf u}_{s})\}$ be a minimal system of binomial generators of $I_{A}$ up to radical, where $s\leq 2r-2$, and let $C=\{{\bf u}_{1},\ldots,{\bf u}_{s}\}$. Let $C_{1}=\{{\bf u}_{1},\ldots,{\bf u}_{k}\}$ and $C_{2}=\{{\bf u}_{k+1},\ldots,{\bf u}_{s}\}$, where $k=\lfloor\frac{s}{2}\rfloor$. Here, $\lfloor y\rfloor$ denotes the greatest integer less than or equal to $y$. Then, $C=C_{1}\cup C_{2}$. Since the dimension of ${\rm span}_{\mathbb{Q}}(C_{1})$ is less than or equal to $k$ and $k\leq\frac{s}{2}$, we conclude that the dimension of ${\rm span}_{\mathbb{Q}}(C_{1})$ is less than or equal to $\frac{s}{2}$. Therefore, the dimension of ${\rm span}_{\mathbb{Q}}(C_{1})$ is less than or equal to $r-1$ because $s\leq 2r-2$. But the dimension of ${\rm ker}_{\mathbb{Q}}(A)$ is equal to $r$, so the dimension of ${\rm span}_{\mathbb{Q}}(C_{1})$ is strictly smaller than the dimension of ${\rm ker}_{\mathbb{Q}}(A)$. Thus, ${\rm span}_{\mathbb{Q}}(C_{1})\subsetneqq{\rm ker}_{\mathbb{Q}}(A)$. Note that $\frac{s}{2}<\lfloor\frac{s}{2}\rfloor+1$, and therefore, $\frac{s}{2}+1>s-\lfloor\frac{s}{2}\rfloor$. But $2r-2\geq s$, so $r\geq\frac{s}{2}+1$ and therefore $r>s-\lfloor\frac{s}{2}\rfloor$. Since the dimension of ${\rm span}_{\mathbb{Q}}(C_{2})$ is less than or equal to $s-\lfloor\frac{s}{2}\rfloor$ and the dimension of ${\rm ker}_{\mathbb{Q}}(A)$ is equal to $r$, we have that the dimension of ${\rm span}_{\mathbb{Q}}(C_{2})$ is strictly smaller than the dimension of ${\rm ker}_{\mathbb{Q}}(A)$. Consequently, ${\rm span}_{\mathbb{Q}}(C_{2})\subsetneqq{\rm ker}_{\mathbb{Q}}(A)$. By Corollary 2.2, the ideal $I_{A}$ is radical splittable. $\square$
Corollary 2.6.
Let $K$ be a field of any characteristic and let $I_{A}$ be a toric ideal of height $r\geq 3$. If ${\rm bar}(I_{A})=r+1$, then $I_{A}$ is radical splittable.
Proof. Since $r\geq 3$, we have that $r+1\leq 2r-2$. By Theorem 2.5, the ideal $I_{A}$ is radical splittable. $\square$
Let $d_{1},\ldots,d_{n}$ be positive integers and $a_{i,j}$ be integers, where $1\leq i\leq r$, $1\leq j\leq n$ and for all $i=1,\ldots,r$ at least one of $a_{i,1},\ldots,a_{i,n}$ is nonzero. Let $A$ be the set of columns of the matrix
$\left(\begin{array}[]{ccccccc}d_{1}&0&\cdots&0&a_{1,1}&\cdots&a_{r,1}\\ 0&d_{2}&\cdots&0&a_{1,2}&\cdots&a_{r,2}\\ \cdots&\cdots&\cdots&\cdots&\cdots&\cdots&\cdots\\ 0&0&\cdots&d_{n}&a_{1,n}&\cdots&a_{r,n}\end{array}\right),$
and let $I_{A}\subset K[x_{1},\ldots,x_{n+r}]$ be the corresponding toric ideal of height $r$. The toric variety $V(I_{A})\subset K^{n+r}$ is called simplicial if $a_{i,j}$ is nonnegative for every $1\leq i\leq r$ and $1\leq j\leq n$. We say that a simplicial toric variety $V(I_{A})\subset K^{n+r}$ has a full parametrization if $a_{i,j}\neq 0$ for all $(i,j)$.
Remark 2.7.
Let $K$ be a field of positive characteristic. Given a simplicial toric variety $V(I_{A})$ with full parametrization, we have from [1, Theorem 1] that $I_{A}$ is a set-theoretic complete intersection on binomials, and therefore $I_{A}$ is radical splittable by Proposition 2.3.
The following proposition demonstrates that the toric ideal $I_{A}$, associated with a simplicial toric variety $V(I_{A})$ that admits full parametrization, is radical splittable whenever ${\rm ht}(I_{A})\geq 3$.
Proposition 2.8.
Let $K$ be a field of any characteristic and let $V(I_{A})\subset K^{n+r}$ be a simplicial toric variety with full parametrization, where ${\rm ht}(I_{A})=r\geq 3$. Then $I_{A}$ is radical splittable.
Proof. If ${\rm char}(K)>0$, then $I_{A}$ is radical splittable by Remark 2.7. Suppose that ${\rm char}(K)=0$. If $I_{A}$ is a complete intersection, that is $\mu(I_{A})={\rm ht}(I_{A})$, then $I_{A}$ is splittable by [12, Corollary 3.3], and therefore $I_{A}$ is radical splittable. Suppose that $I_{A}$ is not a complete intersection. Then ${\rm bar}(I_{A})=r+1$ by [1, Theorem 2], and therefore $I_{A}$ is radical splittable by Corollary 2.6. $\square$
Remark 2.9.
Let $K$ be a field of any characteristic and let $A=\{a_{1},\ldots,a_{n}\}$ be a set of distinct positive integers, where $n\geq 4$. Then $V(I_{A})\subset K^{n}$ is a simplicial toric variety with full parametrization and also $I_{A}$ has height $n-1$. By Proposition 2.8, the toric ideal $I_{A}$ is radical splittable.
Next, we provide an example of a simplicial toric variety that does not admit a full parametrization, such that $I_{A}$ is radical splittable.
Example 2.10.
Let $A$ be the set of columns of the matrix
$\left(\begin{array}[]{cccccccccc}3&0&0&2&1&0&0&2&1&1\\ 0&3&0&1&2&2&1&0&0&1\\ 0&0&3&0&0&1&2&1&2&1\end{array}\right).$
Then $I_{A}$ is a toric ideal of height 7 which is generated up to radical by the binomials $B({\bf u}_{i})$, $1\leq i\leq 11$, where
$\begin{split}{\bf u}_{1}=(2,1,0,-3,0,0,0,0,0,0),\\ {\bf u}_{2}=(1,0,0,-2,1,0,0,0,0,0),\\ {\bf u}_{3}=(1,2,0,0,-3,0,0,0,0,0),\\ {\bf u}_{4}=(0,2,1,0,0,-3,0,0,0,0),\\ {\bf u}_{5}=(0,1,0,0,0,-2,1,0,0,0),\\ {\bf u}_{6}=(0,1,2,0,0,0,-3,0,0,0),\\ {\bf u}_{7}=(2,0,1,0,0,0,0,-3,0,0),\\ {\bf u}_{8}=(1,0,0,0,0,0,0,-2,1,0),\\ {\bf u}_{9}=(1,0,2,0,0,0,0,0,-3,0),\\ {\bf u}_{10}=(1,0,0,-1,0,0,0,-1,0,1),\\ \text{and}\ {\bf u}_{11}=(0,0,0,1,0,0,1,0,0,-2).\end{split}$
From Example 3.7 in [9], we know that if the characteristic of the field $K$ is equal to 0, then ${\rm bar}(I_{A})=11$. In this case, we may take $C=\{{\bf u}_{1},\ldots,{\bf u}_{11}\}$, $C_{1}=\{{\bf u}_{1},{\bf u}_{3},{\bf u}_{4},{\bf u}_{6},{\bf u}_{7},{\bf u}_{9}\}$, and $C_{2}=\{{\bf u}_{2},{\bf u}_{5},{\bf u}_{8},{\bf u}_{10},{\bf u}_{11}\}$. We have that ${\rm span}_{\mathbb{Q}}(C_{1})\subsetneqq{\rm ker}_{\mathbb{Q}}(A)$ and ${\rm span}_{\mathbb{Q}}(C_{2})\subsetneqq{\rm ker}_{\mathbb{Q}}(A)$, since dim$({\rm ker}_{\mathbb{Q}}(A))=7.$ From Corollary 2.2 the ideal $I_{A}$ is radical splittable when the characteristic of $K$ equals 0. Next, we apply the technique described in the proof of Theorem 2.1 to find sets $A_{1}$ and $A_{2}$ such that ${\rm span}_{\mathbb{Q}}(C_{i})={\rm ker}_{\mathbb{Q}}(A_{i})$ for $1\leq i\leq 2$. In this way, we obtain the following two matrices:
$\left(\begin{array}[]{cccccccccc}3&0&0&2&1&0&0&2&1&0\\ 0&3&0&1&2&2&1&0&0&0\\ 0&0&3&0&0&1&2&1&2&0\\ 0&0&0&0&0&0&0&0&0&1\end{array}\right),$
$\left(\begin{array}[]{cccccccccc}2&1&0&1&0&0&-1&1&0&0\\ -1&-1&0&-1&-1&0&1&0&1&0\\ 0&-1&0&1&2&0&1&0&0&1\\ 0&0&1&0&0&0&0&0&0&0\\ 0&2&0&0&0&1&0&0&0&0\end{array}\right).$
Let $A_{1}$ be the set of columns of the first matrix and $A_{2}$ the set of columns of the second matrix. Then $I_{A}={\rm rad}(I_{A_{1}}+I_{A_{2}})$.
If $K$ is a field of characteristic $3$, then ${\rm bar}(I_{A})=7$ by [9, Example 3.7]; namely, $I_{A}$ is a set-theoretic complete intersection on binomials, and therefore $I_{A}$ is radical splittable by Proposition 2.3. In this case, $I_{A}$ is generated up to radical by the binomials $B({\bf u}_{i})$ for $i\in\{1,3,4,6,7,9\}$, and $B({\bf v})$, where ${\bf v}=(1,1,1,0,0,0,0,0,0,-3)$. Hence, $I_{A}={\rm rad}(I_{A_{1}}+\langle x_{1}x_{2}x_{3}-x_{10}^{3}\rangle)$. If $K$ is a field of positive characteristic $p\neq 3$, then, by [9, Example 3.7], $I_{A}$ is generated up to radical by the 8 binomials $B({\bf u}_{i})$, $i\in\{1,3,4,6,7,9,10,11\}$, and therefore $I_{A}$ is radical splittable by Proposition 2.3 and Corollary 2.6. In this case, $I_{A}={\rm rad}(I_{A_{1}}+I_{A_{3}})$ where $I_{A_{3}}=\langle x_{1}x_{10}-x_{4}x_{8},x_{4}x_{7}-x_{10}^{2},x_{1}x_{7}-x_{8}x_{10}\rangle$.
3. Applications to toric ideals of bipartite graphs
Our aim in Section 3 is to prove that if $G$ is a bipartite graph, then ${\rm Split}_{\rm rad}(I_{G})={\rm Split}(I_{G})$, and that both invariants are equal to $3$ when $G$ is a complete bipartite graph.
The radical generation of a toric ideal is a particularly difficult problem. In the literature, certain graphs or simplicial complexes (see [8, 10]) provide crucial information for identifying binomials in a minimal generating set of a toric ideal $I_{A}$ up to radical, or for computing the binomial arithmetical rank of $I_{A}$. In this section and the next, we repeatedly use techniques developed in [8].
First, we recall some material from [8]. The support of a monomial ${\bf x}^{\bf u}$ of $K[x_{1},\ldots,x_{n}]$ is ${\rm supp}({\bf x}^{\bf u})=\{i\mid x_{i}\ \textrm{divides}\ {\bf x}^{\bf u}\}$. A binomial $B={\bf x}^{\bf u}-{\bf x}^{\bf v}$ is called indispensable of $I_{A}$ if every system of binomial generators of $I_{A}$ contains $B$ or $-B$, while a monomial ${\bf x}^{\bf u}$ is called indispensable of $I_{A}$ if every system of binomial generators of $I_{A}$ contains a binomial $B$ in which ${\bf x}^{\bf u}$ appears as one of the monomials. Let $\mathcal{T}$ be the set of all $E\subset\{1,\ldots,n\}$ for which there exists an indispensable monomial $M$ of $I_{A}$ such that $E={\rm supp}(M)$. Let $\mathcal{T}_{\rm min}$ denote the set of minimal elements of $\mathcal{T}$. We associate to $I_{A}$ a graph $\Gamma_{A}$ with vertices the elements of $\mathcal{T}_{\rm min}$. Two distinct vertices $E,E^{\prime}$ are joined by an edge if there exists a binomial $B({\bf u})\in I_{A}$ such that $E={\rm supp}({\bf x}^{{\bf u}^{+}})$ and $E^{\prime}={\rm supp}({\bf x}^{{\bf u}^{-}})$.
A set $M=\{T_{1},\ldots,T_{s}\}$ is called a $\{0,1\}$-matching in $\Gamma_{A}$ if every $T_{k},1\leq k\leq s$ is either an edge of $\Gamma_{A}$ or a singleton set of the form $\{E\}$, where $E$ is a vertex of $\Gamma_{A}$, and also $T_{k}\cap T_{l}=\emptyset$ for all distinct indices $k,l$. Given a $\{0,1\}$-matching $M=\{T_{1},\ldots,T_{s}\}$ in $\Gamma_{A}$, we let ${\rm supp}(\mathcal{M})=\cup_{i=1}^{s}T_{i}$ which is a subset of the vertices $\mathcal{T}_{\rm min}$. A $\{0,1\}$-matching $\mathcal{M}$ in $\Gamma_{A}$ is called a maximal $\{0,1\}$-matching if ${\rm supp}(\mathcal{M})$ has the maximum possible cardinality among all $\{0,1\}$-matchings. Given a maximal $\{0,1\}$-matching =$\mathcal{M}=\{T_{1},\ldots,T_{s}\}$ in $\Gamma_{A}$, we write ${\rm card}(\mathcal{M})$ for the cardinality $s$ of the set $\mathcal{M}$. By $\delta(\Gamma_{A})_{\{0,1\}}$, we denote the minimum of the set
$\{\textrm{card}(N)\mid N\ \textrm{is a maximal}\{0,1\}-\textrm{matching in}\ \Gamma_{A}\}.$
By [8, Theorem 2.6], it holds that ${\rm bar}(I_{A})\geq\delta(\Gamma_{A})_{\{0,1\}}$.
The rest of this section is devoted to studying our problem in the case of toric ideals of graphs, particularly those arising from bipartite graphs.
Let $G$ be a finite, simple, connected graph in the vertex set $\{v_{1},\ldots,v_{n}\}$ with the edge set $E(G)=\{e_{1},\ldots,e_{m}\}$. For each edge $e=\{v_{i},v_{j}\}$ of $G$, we associate a vector ${\bf a}_{e}\in\{0,1\}^{n}$ defined as follows: the $i$th entry of ${\bf a}_{e}$ is $1$, the $j$th entry is $1$, and all other entries are zero. We denote the toric ideal $I_{A_{G}}$ in $K[e_{1},\ldots,e_{m}]$ by $I_{G}$, where $A_{G}=\{{\bf a}_{e}\mid e\in E(G)\}\subset\mathbb{N}^{n}$.
Given an even closed walk $\gamma=(e_{i_{1}},e_{i_{2}},\ldots,e_{i_{2q}})$ in $G$, we define the binomial $f_{\gamma}=\prod_{k=1}^{q}e_{i_{2k-1}}-\prod_{k=1}^{q}e_{i_{2k}}\in I_{G}$. By [17, Proposition 10.1.5], the ideal $I_{G}$ is generated by all the binomials $f_{\gamma}$, where $\gamma$ is an even closed walk of $G$. Note that if $G$ is a bipartite graph, then it contains no odd cycles, and thus $I_{G}$ is generated by the binomials $f_{\gamma}$, where $\gamma$ is an even cycle in $G$.
We grade $K[e_{1},\ldots,e_{m}]$ by setting ${\rm deg}_{A_{G}}(x_{i})={\bf a}_{e_{i}}$ for every $1\leq i\leq m$. Given a binomial $B({\bf u})\in I_{G}$, where ${\bf u}={\bf u}^{+}-{\bf u}^{-}\in{\rm ker}_{\mathbb{Z}}(A_{G})$, we have
${\rm deg}_{A_{G}}({\bf e}^{{\bf u}^{+}})={\rm deg}_{A_{G}}({\bf e}^{{\bf u}^{-}}).$
For simplicity, we shall write ${\rm ker}_{\mathbb{Z}}(G)$ instead of ${\rm ker}_{\mathbb{Z}}(A_{G})$ and ${\rm ker}_{\mathbb{Q}}(G)$ instead of ${\rm ker}_{\mathbb{Q}}(A_{G})$.
In [11, Theorem 3.2], it is proved that the toric ideal of $K_{n}$ is subgraph splittable for all $n\geq 4$, where $K_{n}$ denotes the complete graph with $n$ vertices. Note that for $n<4$, the ideal $I_{K_{n}}$ is the zero ideal. Recall that $I_{K_{n}}$ is subgraph splittable if there exist subgraphs $G_{1}$ and $G_{2}$ of $K_{n}$ such that $I_{K_{n}}=I_{G_{1}}+I_{G_{2}}.$ In terms of splitting numbers, this means that both ${\rm Split}(I_{K_{n}}),{\rm Split}_{\rm rad}(I_{K_{n}})$ are equal to two for $n\geq 4.$ In contrast, [12, Theorem 4.5], shows that the toric ideal of the complete bipartite graph $K_{m,n}$ is not splittable. Thus, ${\rm Split}(I_{K_{m,n}})$ is not equal to two. The aim of the next two theorems is to prove that both the splitting number and the radical splitting number of the toric ideal of a complete bipartite graph $K_{m,n}$ are equal to three, where $m,n\geq 2$ and $(m,n)\not=(2,2)$. Moreover, we prove that $I_{K_{m,n}}=I_{G_{1}}+I_{G_{2}}+I_{G_{3}}$, for three specific subgraphs of $K_{m,n}$.
Theorem 3.1.
Let $G$ be a bipartite graph, then ${\rm Split}(I_{G})={\rm Split}_{\rm rad}(I_{G})$.
Proof. Let $V_{1}=\{x_{1},\ldots,x_{m}\}$ and $V_{2}=\{y_{1},\ldots,y_{n}\}$ be a bipartition of the graph $G$, such that every edge of $G$ connects a vertex in $V_{1}$ to a vertex in $V_{2}$. We denote the edge $\{x_{i},y_{j}\}$ by $b_{ij}$.
Suppose that ${\rm Split}_{\rm rad}(I_{G})=r$, then there exists a minimal system of binomial generators $\{B({\bf u})\mid{\bf u}\in C\subset{\rm ker}_{\mathbb{Z}}(G)\}$ of the toric ideal $I_{G}$ up to radical, and sets $C_{i}$, $1\leq i\leq r$ such that $C=\cup_{i=1}^{r}C_{i}$ and ${\rm span}_{\mathbb{Q}}(C_{i})\subsetneqq{\rm ker}_{\mathbb{Q}}(G)$ for every $1\leq i\leq r$. By [15, Theorem 2.3], the ideal $I_{G}$ has a unique minimal system of binomial generators $S$ consisting of all binomials of the form $f_{\gamma}$, where $\gamma$ is an even cycle of $G$ with no chord. We will show that each minimal generator $f_{\gamma}\in S$ belongs to the set $\{B({\bf u})\mid{\bf u}\in C\subset{\rm ker}_{\mathbb{Z}}(G)\}$. Let $\gamma=(x_{i_{1}},y_{j_{1}},x_{i_{2}},y_{j_{2}},\ldots,x_{i_{s}},y_{j_{s}})$ be the cycle and $f_{\gamma}={\bf e}^{{\bf v}^{+}}-{\bf e}^{{\bf v}^{-}}=b_{i_{1}j_{1}}b_{i_{2}j_{2}}\cdots b_{i_{s}j_{s}}-b_{i_{1}j_{s}}b_{i_{2}j_{1}}\cdots b_{i_{s}j_{s-1}}$, where ${\bf v}={\bf v}^{+}-{\bf v}^{-}\in{\rm ker}_{\mathbb{Z}}(G)$. The cycle $\gamma$ has no chords; thus, the induced subgraph $W$ of $G$ on the vertex set $\{x_{i_{1}},x_{i_{2}},\ldots,x_{i_{s}},y_{j_{1}},y_{j_{2}},\ldots,y_{j_{s}}\}$ has edge set $E(W)=\{b_{i_{1}j_{1}},b_{i_{2}j_{1}},b_{i_{2}j_{2}},b_{i_{3}j_{2}},\ldots,b_{i_{s}j_{s}},b_{i_{1}j_{s}}\}$. Since $f_{\gamma}$ is an indispensable binomial of $I_{G}$, both ${\bf e}^{{\bf v}^{+}}$ and ${\bf e}^{{\bf v}^{-}}$ are indispensable monomials; thus, we have $E={\rm supp}({\bf e}^{{\bf v}^{+}})\in\mathcal{T}_{\rm min}$. By [8, Theorem 2.4], there exists a vector ${\bf u}\in C$ such that ${\rm supp}({\bf e}^{{\bf u}^{+}})=E$ or ${\rm supp}({\bf e}^{{\bf u}^{-}})=E$. Without loss of generality, we assume that ${\rm supp}({\bf e}^{{\bf u}^{+}})=E$, and also that $B({\bf u})={\bf e}^{{\bf u}^{+}}-{\bf e}^{{\bf u}^{-}}$ is irreducible. Let ${\bf e}^{{\bf u}^{+}}=b_{i_{1}j_{1}}^{a_{1}}b_{i_{2}j_{2}}^{a_{2}}\cdots b_{i_{s}j_{s}}^{a_{s}}$. Then ${\rm deg}_{A_{G}}({\bf e}^{{\bf u}^{+}})=\Sigma_{l\in[s]}a_{l}\varepsilon_{i_{l}}+\Sigma_{l\in[s]}a_{l}\varepsilon_{m+j_{l}}$, where $\varepsilon_{t}$ is the $t$-th unit vector of $\mathbb{Z}^{m+n}$. Since ${\rm deg}_{A_{G}}({\bf e}^{{\bf u}^{+}})={\rm deg}_{A_{G}}({\bf e}^{{\bf u}^{-}})$, the variables that appear in ${\bf e}^{{\bf u}^{-}}$ correspond to edges of $W$. Moreover, since ${\bf e}^{{\bf u}^{+}}$ and ${\bf e}^{{\bf u}^{-}}$ have no common factor, the variables corresponding to the edges $b_{i_{1}j_{1}},b_{i_{2}j_{2}},\ldots,b_{i_{s}j_{s}}$ cannot appear in ${\bf e}^{{\bf u}^{-}}$. Thus, ${\bf e}^{{\bf u}^{-}}$ must be of the form $b_{i_{1}j_{s}}^{c_{1}}b_{i_{2}j_{1}}^{c_{2}}\cdots b_{i_{s}j_{s-1}}^{c_{s}}$. Then, from ${\rm deg}_{A_{G}}({\bf e}^{{\bf u}^{+}})={\rm deg}_{A_{G}}({\bf e}^{{\bf u}^{-}})$, we have
$\sum_{l\in[s]}a_{l}\varepsilon_{i_{l}}+\sum_{l\in[s]}a_{l}\varepsilon_{m+j_{l}}=\sum_{l\in[s]}c_{l}\varepsilon_{i_{l}}+c_{1}\varepsilon_{m+{j_{s}}}+\sum_{\begin{subarray}{c}l\in[s]\\ l\neq 1\end{subarray}}c_{l}\varepsilon_{m+j_{l-1}},$
and therefore $a_{1}=a_{2}=\cdots=a_{s}=c_{1}=c_{2}=\cdots=c_{s}$. Moreover, since $B({\bf u})$ is irreducible, we have $a_{1}=a_{2}=\cdots=a_{s}=c_{1}=c_{2}=\cdots=c_{s}=1$. Therefore,
$B({\bf u})=b_{i_{1}j_{1}}b_{i_{2}j_{2}}\cdots b_{i_{s}j_{s}}-b_{i_{1}j_{s}}b_{i_{2}j_{1}}\cdots b_{i_{s}j_{s-1}}.$
Consequently, $f_{\gamma}=B({\bf u})$. Thus, $S\subset\{B({\bf u})\mid{\bf u}\in C\subset{\rm ker}_{\mathbb{Z}}(G)\}$. In fact $S=\{B({\bf u})\mid{\bf u}\in C\subset{\rm ker}_{\mathbb{Z}}(G)\}$, since $\{B({\bf u})\mid{\bf u}\in C\subset{\rm ker}_{\mathbb{Z}}(G)\}$ is a minimal system of generators of $I_{G}$ up to radical. By Theorem 2.1, there exist toric ideals $I_{A_{i}}$ with $I_{A_{i}}\neq I_{G}$ for each $i$ with $1\leq i\leq r$, such that $I_{G}=I_{A_{1}}+\cdots+I_{A_{r}}$. Consequently ${\rm Split}(I_{G})\leq r$ and therefore ${\rm Split}(I_{G})=r$ since always ${\rm Split}_{\rm rad}(I_{G})\leq{\rm Split}(I_{G})$. $\square$
A bipartite graph $G$ is called a complete bipartite graph if its vertex set can be partitioned into two subsets $V_{1}$ and $V_{2}$ such that every vertex of $V_{1}$ is connected to every vertex of $V_{2}$. It is denoted by $K_{m,n}$, where $m$ and $n$ are the numbers of vertices in $V_{1}$ and $V_{2}$ respectively. The following example provides a splitting of $I_{K_{3,3}}$.
Example 3.2.
Let $K_{3,3}$ be the complete bipartite graph on the vertex set $\{x_{1},x_{2},x_{3}\}\cup\{y_{1},y_{2},y_{3}\}$ with edges $E(K_{3,3})=\{b_{ij}\mid 1\leq i\leq 3,1\leq j\leq 3\}$, where $b_{ij}=\{x_{i},y_{j}\}$. Then, $I_{K_{3,3}}$ is minimally generated by the binomials $b_{21}b_{33}-b_{23}b_{31}$, $b_{22}b_{33}-b_{23}b_{32}$, $b_{21}b_{32}-b_{22}b_{31}$, $b_{11}b_{23}-b_{13}b_{21}$, $b_{12}b_{23}-b_{13}b_{22}$, $b_{11}b_{22}-b_{12}b_{21}$, $b_{11}b_{33}-b_{13}b_{31}$, $b_{12}b_{33}-b_{13}b_{32}$, and $b_{11}b_{32}-b_{12}b_{31}$. Let $G_{1}$ be the induced subgraph of $K_{3,3}$ on the vertex set $\{x_{2},x_{3}\}\cup\{y_{1},y_{2},y_{3}\}$, then $I_{G_{1}}$ is minimally generated by the binomials $b_{21}b_{33}-b_{23}b_{31}$, $b_{22}b_{33}-b_{23}b_{32}$, and $b_{21}b_{32}-b_{22}b_{31}$. Let $G_{2}$ be the induced subgraph of $K_{3,3}$ on the vertex set $\{x_{1},x_{2}\}\cup\{y_{1},y_{2},y_{3}\}$, then $I_{G_{2}}$ is minimally generated by the binomials $b_{11}b_{23}-b_{13}b_{21}$, $b_{12}b_{23}-b_{13}b_{22}$, and $b_{11}b_{22}-b_{12}b_{21}$. Let $G_{3}$ be the induced subgraph of $K_{3,3}$ on the vertex set $\{x_{1},x_{3}\}\cup\{y_{1},y_{2},y_{3}\}$, then $I_{G_{3}}$ is minimally generated by the binomials $b_{11}b_{33}-b_{13}b_{31}$, $b_{12}b_{33}-b_{13}b_{32}$, and $b_{11}b_{32}-b_{12}b_{31}$. Thus, $I_{K_{3,3}}=I_{G_{1}}+I_{G_{2}}+I_{G_{3}}$ constitutes a splitting of $I_{K_{3,3}}$.
Next, we explicitly compute ${\rm Split}_{\rm rad}(I_{K_{m,n}})$.
Theorem 3.3.
The splitting number and the radical splitting number of the toric ideal of a complete bipartite graph $K_{m,n}$ are both equal to three, where $m,n\geq 2$ and $(m,n)\not=(2,2)$.
Proof. Let $V_{1}=\{x_{1},\ldots,x_{m}\}$, $V_{2}=\{y_{1},\ldots,y_{n}\}$ be the bipartition of the complete bipartite graph $K_{m,n}$ and $E(K_{m,n})=\{b_{ij}|1\leq i\leq m,1\leq j\leq n\}$, where $b_{ij}=\{x_{i},y_{j}\}$. By [12, Theorem 4.5], the ideal $I_{K_{m,n}}$ is not splittable, thus ${\rm Split}(I_{K_{m,n}})\neq 2$, and therefore ${\rm Split}(I_{K_{m,n}})\geq 3$. The ideal $I_{K_{m,n}}$ is minimally generated by the $2\times 2$ minors of the matrix $M=(b_{ij})$, see [17, Proposition 10.6.2]. Thus, $I_{K_{m,n}}$ is minimally generated by all binomials of the form $b_{ij}b_{kl}-b_{il}b_{kj}$, such that $(b_{ij},b_{kj},b_{kl},b_{il})$ is a cycle of length 4 of $K_{m,n}$. By the hypothesis of the theorem, at least one of the sets $V_{1}$, $V_{2}$ contains at least three vertices. Suppose $V_{1}$ contains three or more vertices. Consider the following three induced subgraphs of $K_{m,n}$: the graph $G_{1}$, with vertex set $(V_{1}\setminus\{x_{1}\})\cup V_{2}$; the graph $G_{2}$, with vertex set $(V_{1}\setminus\{x_{m}\})\cup V_{2}$; and the graph $G_{3}$, with vertex set $\{x_{1},x_{m}\}\cup V_{2}$. Let $f_{w}=b_{ij}b_{kl}-b_{il}b_{kj}$ be a minimal generator of the toric ideal of $K_{m,n}$, where $b_{ij}=\{x_{i},y_{j}\}$, $b_{kj}=\{x_{k},y_{j}\}$, $b_{kl}=\{x_{k},y_{l}\}$, and $b_{il}=\{x_{i},y_{l}\}$. Notice that $x_{i},x_{k}\in V_{1}$ and $y_{j},y_{l}\in V_{2}$. We distinguish the following cases.
  1. (1)
    If $i\neq 1$, then $f_{w}$ belongs to $I_{G_{1}}$.
  2. (2)
    If $i=1$ and $k\neq m$, then $f_{w}$ belongs to $I_{G_{2}}$.
  3. (3)
    If $i=1$ and $k=m$, then $f_{w}$ belongs to $I_{G_{3}}$.
So, every minimal generator of $I_{K_{m,n}}$ belongs to at least one of the toric ideals $I_{G_{1}}$, $I_{G_{2}}$, or $I_{G_{3}}$, and therefore $I_{K_{m,n}}\subseteq I_{G_{1}}+I_{G_{2}}+I_{G_{3}}$. Clearly, $I_{G_{i}}\neq I_{K_{m,n}}$ for all $1\leq i\leq 3$. Thus, $I_{K_{m,n}}=I_{G_{1}}+I_{G_{2}}+I_{G_{3}}$, and ${\rm Split}(I_{K_{m,n}})=3$. Since $K_{m,n}$ is bipartite, it follows from Theorem 3.1 that ${\rm Split}_{\rm rad}(I_{K_{m,n}})={\rm Split}(I_{K_{m,n}})$. $\square$
4. The case of height 2 toric ideals
In this section, we study the special case in which $I_{A}\subset K[x_{1},\ldots,x_{n}]$ is a toric ideal of height $2$. The next theorem shows that ${\rm Split}(I_{A})=\mu(I_{A})$ and ${\rm Split}_{\rm rad}(I_{A})={\rm bar}(I_{A})$ whenever ${\rm ht}(I_{A})=2$.
Theorem 4.1.
Let $I_{A}$ be a toric ideal of height 2. Then the splitting number of $I_{A}$ (respectively, the radical splitting number of $I_{A}$) equals the minimal number of generators $\mu(I_{A})$ of $I_{A}$ (respectively, the binomial arithmetical rank of $I_{A}$).
Proof. Suppose that ${\rm Split}(I_{A})=r$ (respectively, ${\rm Split}_{\rm rad}(I_{A})=r$), and let $I_{A}=I_{A_{1}}+\cdots+I_{A_{r}}$ be a splitting of $I_{A}$ (respectively, let $I_{A}={\rm rad}(I_{A_{1}}+\cdots+I_{A_{r}})$ be a radical splitting of $I_{A}$). Since $I_{A_{i}}\subsetneqq I_{A}$, we have that ${\rm ht}(I_{A_{i}})<{\rm ht}(I_{A})$, and therefore ${\rm ht}(I_{A_{i}})=1$ for every $1\leq i\leq r$. Thus, each $I_{A_{i}}$ is a principal ideal, and let $T_{i}$ be a minimal binomial generating set of $I_{A_{i}}$, which is a singleton. Then $D=\cup_{i=1}^{r}T_{i}$ generates $I_{A}$ (respectively, $D$ generates $I_{A}$ up to radical). So $r\geq\mu(I_{A})$ (respectively, $r\geq{\rm bar}(I_{A})$), and therefore $\mu(I_{A})=r$ (respectively, ${\rm bar}(I_{A})=r$), since ${\rm Split}(I_{A})\leq\mu(I_{A})$ (respectively, ${\rm Split}_{\rm rad}(I_{A})\leq{\rm bar}(I_{A})$). $\square$
The next corollary provides a necessary and sufficient condition for $I_{A}$ to be radical splittable.
Corollary 4.2.
Let $K$ be a field of any characteristic and let $I_{A}$ be a toric ideal of height 2. Then $I_{A}$ is radical splittable if and only if $I_{A}$ is a set-theoretic complete intersection on binomials.
Proof. Since ${\rm Split}_{\rm rad}(I_{A})={\rm bar}(I_{A})$ by Theorem 4.1, we have that $I_{A}$ is radical splittable if and only if ${\rm bar}(I_{A})=2$; in other words $I_{A}$ is a set-theoretic complete intersection on binomials. $\square$
Remark 4.3.
Suppose that ${\rm char}(K)=0$ and that the height of $I_{A}$ is 2. If $I_{A}$ is not a complete intersection, then $I_{A}$ is not a set-theoretic complete intersection on binomials by [2, Theorem 4], and therefore ${\rm Split}_{\rm rad}(I_{A})\geq 3$.
The next corollary provides bounds for ${\rm Split}_{\rm rad}(I_{A})$ whenever $V(I_{A})$ is a simplicial toric variety.
Corollary 4.4.
Let $K$ be a field of any characteristic, and let $V(I_{A})$ be a simplicial toric variety, where $I_{A}$ is a toric ideal of height 2. Then $2\leq{\rm Split}_{\rm rad}(I_{A})\leq 3$.
Proof. By [1, Theorem 3], we have $2\leq{\rm bar}(I_{A})\leq 3$, and therefore $2\leq{\rm Split}_{\rm rad}(I_{A})\leq 3$. $\square$
Next, we provide an example of a simplicial toric variety $V(I_{A})$ with full parametrization, such that ${\rm ht}(I_{A})=2$ and ${\rm Split}_{\rm rad}(I_{A})=3$ when the characteristic of $K$ is zero.
Example 4.5.
Let $A$ be the set of columns of the matrix
$\left(\begin{array}[]{cccc}2&1&2&0\\ 3&0&2&5\end{array}\right).$
Then $I_{A}$ is a toric ideal of height 2 which is minimally generated by the binomials $x_{1}^{3}-x_{2}^{2}x_{3}^{2}x_{4}$, $x_{3}^{3}-x_{1}^{2}x_{2}^{2}$, and $x_{1}x_{3}-x_{2}^{4}x_{4}$. By Corollary 4.4, it holds $2\leq{\rm Split}_{\rm rad}(I_{A})\leq 3$. If ${\rm char}(K)>0$, then $I_{A}$ is radical splittable by Remark 2.7, and therefore ${\rm Split}_{\rm rad}(I_{A})=2$. If ${\rm char}(K)=0$, then ${\rm Split}_{\rm rad}(I_{A})\geq 3$ by Remark 4.3, and therefore ${\rm Split}_{\rm rad}(I_{A})=3$.
We now provide an example of a non-simplicial toric variety $V(I_{A})$ such that ${\rm ht}(I_{A})=2$ and ${\rm Split}_{\rm rad}(I_{A})=3$, in any characteristic.
Example 4.6.
Let $A$ be the set of columns of the matrix
$\left(\begin{array}[]{ccccc}33&-4&1&0&1\\ 23&-3&0&-9&0\\ -34&6&0&3&1\end{array}\right).$
Then $I_{A}$ is a toric ideal of height 2 which is minimally generated by the binomials $x_{5}^{21}-x_{1}^{3}x_{2}^{20}x_{3}^{2}x_{4}$, $x_{1}^{3}x_{2}^{17}x_{4}^{2}-x_{3}^{25}x_{5}^{6}$, $x_{3}^{52}-x_{1}^{3}x_{2}^{14}x_{4}^{3}x_{5}^{9}$, and $x_{2}^{3}x_{3}^{27}-x_{4}x_{5}^{15}$. We have that $I_{A}={\rm rad}(x_{5}^{21}-x_{1}^{3}x_{2}^{20}x_{3}^{2}x_{4},x_{3}^{52}-x_{1}^{3}x_{2}^{14}x_{4}^{3}x_{5}^{9},x_{1}^{3}x_{2}^{17}x_{4}^{2}-x_{3}^{25}x_{5}^{6})$, since $(x_{2}^{3}x_{3}^{27}-x_{4}x_{5}^{15})^{2}$ belongs to the ideal generated by the binomials $x_{5}^{21}-x_{1}^{3}x_{2}^{20}x_{3}^{2}x_{4}$, $x_{1}^{3}x_{2}^{17}x_{4}^{2}-x_{3}^{25}x_{5}^{6}$, and $x_{3}^{52}-x_{1}^{3}x_{2}^{14}x_{4}^{3}x_{5}^{9}$. Thus ${\rm bar}(I_{A})\leq 3$. Then $\mathcal{T}_{{\rm min}}=\{\{3\},\{5\},\{1,2,4\}\}$ and $\Gamma_{A}$ consists only of the three vertices of $\mathcal{T}_{{\rm min}}$. Thus $\delta(\Gamma_{A})_{\{0,1\}}=3$ and therefore ${\rm bar}(I_{A})\geq 3$. Consequently ${\rm bar}(I_{A})=3$, so the radical splitting number of $I_{A}$ equals 3 in any characteristic.
Next, we provide two classes of toric ideals for which the splitting number can be arbitrarily large, while the radical splitting number remains constant. The first class determines a simplicial toric variety, while the second determines a non-simplicial toric variety.
Let $a<b<d$ be positive integers with ${\rm gcd}(a,b,d)=1$ and $a+b=d$. A symmetric monomial curve in $\mathbb{P}^{3}$ is given parametrically by
$(u^{d},u^{d-a}v^{a},u^{d-b}v^{b},v^{d}).$
Let $S$ be the set of columns of the $2\times 4$ matrix
$\left(\begin{array}[]{cccc}d&d-a&d-b&0\\ 0&a&b&d\end{array}\right).$
Then $I_{S}$ is minimally generated (see [3]) by the $b-a+2$ binomials $x_{1}x_{4}-x_{2}x_{3}$ and $x_{1}^{b-a-i}x_{3}^{a+i}-x_{2}^{b-i}x_{4}^{i}$, for every $0\leq i\leq b-a$. By Theorem 4.1, the splitting number of $I_{S}$ equals $b-a+2$.
Note that $I_{S}=I_{A}$ where $A$ is the set of columns of the $2\times 4$ matrix
$\left(\begin{array}[]{cccc}1&1&1&1\\ 0&a&b&a+b\end{array}\right).$
Remark 4.7.
Let $I_{A}$ be the toric ideal of a symmetric monomial curve in $\mathbb{P}^{3}$. Since $V(I_{S})$ is a simplicial toric variety with full parametrization, it follows from Remark 2.7 that the radical splitting number of $I_{A}$ is $2$ whenever ${\rm char}(K)>0$. If ${\rm char}(K)=0$, then ${\rm Split}_{\rm rad}(I_{A})\geq 3$ by Remark 4.3, and hence ${\rm Split}_{\rm rad}(I_{A})=3$.
The Lawrence lifting $\Lambda(A)$ of $A$ is the set of columns of the $6\times 8$ matrix
$\left(\begin{array}[]{cccccccc}1&1&1&1&0&0&0&0\\ 0&a&b&a+b&0&0&0&0\\ 1&0&0&0&1&0&0&0\\ 0&1&0&0&0&1&0&0\\ 0&0&1&0&0&0&1&0\\ 0&0&0&1&0&0&0&1\end{array}\right).$
The toric ideal $I_{\Lambda(A)}\subset K[x_{1},\ldots,x_{4},y_{1},\ldots,y_{4}]$ is generated by all binomials ${\bf x}^{{\bf u}^{+}}{\bf y}^{{\bf u}^{-}}-{\bf x}^{{\bf u}^{-}}{\bf y}^{{\bf u}^{+}}$ where ${\bf u}={\bf u}^{+}-{\bf u}^{-}\in{\rm ker}_{\mathbb{Z}}(A)$. Our aim is to compute ${\rm Split}_{\rm rad}(I_{\Lambda(A)})$.
Given a toric ideal $I_{A}\subset K[x_{1},\ldots,x_{n}]$, a binomial ${\bf x}^{{\bf u}^{+}}-{\bf x}^{{\bf u}^{-}}\in I_{A}$ is called primitive if there is no other binomial ${\bf x}^{{\bf v}^{+}}-{\bf x}^{{\bf v}^{-}}\in I_{A}$ such that ${\bf x}^{{\bf v}^{+}}$ divides ${\bf x}^{{\bf u}^{+}}$ and ${\bf x}^{{\bf v}^{-}}$ divides ${\bf x}^{{\bf u}^{-}}$. The set of the primitive binomials is finite, it is called the Graver basis of $I_{A}$ and is denoted by ${\rm Gr}(I_{A})$. An irreducible binomial ${\bf x}^{{\bf u}^{+}}-{\bf x}^{{\bf u}^{-}}\in I_{A}$ is called a circuit of $I_{A}$ if there exists no binomial ${\bf x}^{{\bf v}^{+}}-{\bf x}^{{\bf v}^{-}}\in I_{A}$ such that ${\rm supp}({\bf x}^{{\bf v}^{+}})\cup{\rm supp}({\bf x}^{{\bf v}^{-}})\subsetneqq{\rm supp}({\bf x}^{{\bf u}^{+}})\cup{\rm supp}({\bf x}^{{\bf u}^{-}})$. Let $\mathcal{C}$ denote the collection of subsets $E\subset\{1,\ldots,n\}$ such that ${\rm supp}({\bf x}^{{\bf u}^{+}})=E$ or ${\rm supp}({\bf x}^{{\bf u}^{-}})=E$, where $B({\bf u})$ is a circuit of $I_{A}$. Let $\mathcal{C}_{\rm min}$ denote the set of the minimal elements of $\mathcal{C}$. By [8, Proposition 2.2], the set $\mathcal{T}_{\rm min}$ is equal to $\mathcal{C}_{\rm min}$.
Proposition 4.8.
For the Lawrence lifting $\Lambda(A)$ of $A$, the following statements hold:
  1. (1)
    The splitting number of $I_{\Lambda(A)}$ is $a+b+2$.
  2. (2)
    The radical splitting number of $I_{\Lambda(A)}$ is $4$.
Proof. Let ${\bf v}_{i}=(b-i,i-a-b,i,a-i)$ for every $0\leq i\leq a+b$, and let ${\bf w}=(1,-1,-1,1)$. By [14, Lemma 2.2], the Graver basis of $I_{A}$ is
${\rm Gr}(I_{A})=\{B({\bf v}_{i})\mid 0\leq i\leq a+b\}\cup\{B({\bf w})\}.$
Then $I_{\Lambda(A)}$ is minimally generated by the $a+b+2$ binomials $B({\bf w},-{\bf w})={\bf x}^{{\bf w}^{+}}{\bf y}^{{\bf w}^{-}}-{\bf x}^{{\bf w}^{-}}{\bf y}^{{\bf w}^{+}}$ and $B({\bf v}_{i},-{\bf v}_{i})={\bf x}^{{\bf v}_{i}^{+}}{\bf y}^{{\bf v}_{i}^{-}}-{\bf x}^{{\bf v}_{i}^{-}}{\bf y}^{{\bf v}_{i}^{+}}$ for every $0\leq i\leq a+b$.
(1) By Theorem 4.1, the splitting number of $I_{\Lambda(A)}$ is $a+b+2$.
(2) Since
${\rm Gr}(I_{A})=\{B({\bf v}_{i})\mid 0\leq i\leq a+b\}\cup\{B({\bf w})\},$
we have that $I_{A}$ has 4 circuits, namely $x_{2}^{b}-x_{1}^{b-a}x_{3}^{a}$, $x_{2}^{a+b}-x_{1}^{b}x_{4}^{a}$, $x_{3}^{b}-x_{2}^{a}x_{4}^{b-a}$, and $x_{3}^{a+b}-x_{1}^{a}x_{4}^{b}$. So $I_{\Lambda(A)}$ has 4 circuits, namely $x_{2}^{b}y_{1}^{b-a}y_{3}^{a}-x_{1}^{b-a}x_{3}^{a}y_{2}^{b}$, $x_{2}^{a+b}y_{1}^{b}y_{4}^{a}-x_{1}^{b}x_{4}^{a}y_{2}^{a+b}$, $x_{3}^{b}y_{2}^{a}y_{4}^{b-a}-x_{2}^{a}x_{4}^{b-a}y_{3}^{b}$, and $x_{3}^{a+b}y_{1}^{a}y_{4}^{b}-x_{1}^{a}x_{4}^{b}y_{3}^{a+b}$. By [5, Proposition 8.7], $I_{\Lambda(A)}$ is generated up to radical by its circuits. Thus ${\rm bar}(I_{\Lambda(A)})\leq 4$. Let $E_{1}={\rm supp}(x_{2}^{b}y_{1}^{b-a}y_{3}^{a})$, $E_{2}={\rm supp}(x_{2}^{a+b}y_{1}^{b}y_{4}^{a})$, $E_{3}={\rm supp}(x_{3}^{b}y_{2}^{a}y_{4}^{b-a})$, $E_{4}={\rm supp}(x_{3}^{a+b}y_{1}^{a}y_{4}^{b})$, $E_{5}={\rm supp}(x_{1}^{b-a}x_{3}^{a}y_{2}^{b})$, $E_{6}={\rm supp}(x_{1}^{b}x_{4}^{a}y_{2}^{a+b})$, $E_{7}={\rm supp}(x_{2}^{a}x_{4}^{b-a}y_{3}^{b})$, and $E_{8}={\rm supp}(x_{1}^{a}x_{4}^{b}y_{3}^{a+b})$. Clearly, $\mathcal{C}_{\rm min}=\{E_{i}\mid 1\leq i\leq 8\}$. Then the graph $\Gamma_{\Lambda(A)}$ consists of four edges, namely $\{E_{1},E_{2}\}$, $\{E_{3},E_{4}\}$, $\{E_{5},E_{6}\}$, and $\{E_{7},E_{8}\}$. Thus, $\delta(\Gamma_{\Lambda(A)})_{\{0,1\}}=4$, and therefore ${\rm bar}(I_{\Lambda(A)})\geq 4$. Consequently, ${\rm bar}(I_{\Lambda(A)})=4$, and thus ${\rm Split}_{\rm rad}(I_{\Lambda(A)})=4$. $\square$
We conclude the paper by presenting a class of toric ideals $I_{A}$ of height $2$ for which both the splitting number and the radical splitting number attain arbitrarily large values. We emphasize that cyclic configurations, known for their extremal properties (see, e.g., [13]), play a key role in our construction.
Given an integer $t$, define
${\bf a}_{t}=\left(\begin{array}[]{c}1\\ t\\ t^{2}\\ \vdots\\ t^{2d-2}\end{array}\right)\in\mathbb{Z}^{2d-1},$
where $d\geq 2$ is an integer. Let $A$ denote the cyclic configuration formed by the columns of the $(2d-1)\times(2d+1)$ Vandermonde matrix
$\left(\begin{array}[]{cccc}{\bf a}_{t_{1}}&{\bf a}_{t_{2}}&\ldots&{\bf a}_{t_{2d+1}}\end{array}\right),$
where $t_{1}<t_{2}<\cdots<t_{2d+1}$ are integers. Any subset of $A$ consisting of $12d-$ 1 vectors is linearly independent, so the dimension of ${\rm ker}_{\mathbb{Q}}(A)$ equals $2$, and therefore ${\rm ht}(I_{A})=2$.
Proposition 4.9.
Let $I_{A}$ be the toric ideal of a cyclic configuration of height two. The splitting number and the radical splitting number of $I_{A}$ can attain arbitrarily large values.
Proof. Any subset of $A$ consisting of $2d$ vectors defines a circuit of $I_{A}$. Consequently, there are exactly $12d+$ 1 circuits in $I_{A}$. By [5, Proposition 8.7], $I_{A}$ is generated up to radical by its circuits, thus ${\rm bar}(I_{A})\leq 2d+1$. Consider an increasing sequence $t_{i_{1}}<t_{i_{2}}<\cdots<t_{i_{2d}}$ obtained by omitting a single element from $\{t_{1},\ldots,t_{2d+1}\}$. Then, by Cramer’s rule, we obtain
$\sum_{j=1}^{2d}(-1)^{j}\det({\bf a}_{t_{i_{1}}},{\bf a}_{t_{i_{2}}},\ldots,{\bf a}_{t_{i_{j-1}}},{\bf a}_{t_{i_{j+1}}},\ldots,{\bf a}_{t_{i_{2d}}})\varepsilon_{i_{j}}$
is a scalar multiple of a circuit, where $\varepsilon_{i}$ denotes the $i$-th unit vector of $\mathbb{Z}^{2d+1}$. Recall that the determinant $\det({\bf a}_{t_{i_{1}}},{\bf a}_{t_{i_{2}}},\ldots,{\bf a}_{t_{i_{j-1}}},{\bf a}_{t_{i_{j+1}}},\ldots,{\bf a}_{t_{i_{2d}}})$ is given by
$\prod_{1\leq k<l\leq 2d}^{k,l\not=j}(t_{i_{l}}-t_{i_{k}})>0.$
Thus the signs of the nonzero components of a circuit alternate. Then every element of $\mathcal{C}_{\rm min}$ has exactly $d$ elements and can be represented as $\{k_{1},k_{2},\ldots,k_{d}\}$, where $11\leq k_{1}<k_{2}<\cdots<k_{d}\leq 2d+$ 1 and $k_{j+1}-k_{j}=2$ for all but at most one index $j$, for which $k_{j+1}-k_{j}=3$. Each circuit contributes two elements in $\mathcal{C}_{\rm min}$, and each element of $\mathcal{C}_{\rm min}$ appears in exactly two circuits. Therefore, the graph $\Gamma_{A}$ is a cycle with $12d+$ 1 vertices and $12d+$ 1 edges. Thus, $\delta(\Gamma_{A})_{\{0,1\}}=d+1$, and therefore ${\rm bar}(I_{A})\geq d+1$. Consequently, $d+1\leq{\rm Split}_{\rm rad}(I_{A})\leq 2d+1$, since ${\rm Split}_{\rm rad}(I_{A})={\rm bar}(I_{A})$ by Theorem 4.1. Since ${\rm Split}_{\rm rad}(I_{A})\leq{\rm Split}(I_{A})$, we conclude that $d+1\leq{\rm Split}(I_{A})$. $\square$